the sound level is 103 dB when one jet engine is roaring 3 more identical jet engines start roaring what is the sound level increased to
Answer #1

Use equation,

β = (10dB)*log(I/I0)

β1 = (10dB)*log(I1/I0)

β3 = (10dB)*log(I3/I0)

β3 - β1 = (10dB)*[log(I3/I0)-log(I1/I0)]

β3 - β1 = (10dB)*[log(I3/I1)]

We are given , I3/I1=3, β1 = 103 dB

β3 - β1 = (10dB)*log3

β3 – 103dB = (10dB)*0.477

β3 – 103dB = 4.77dB

β3 = 4.77dB +103dB

β3 = 107.77 dB